OpenStudy (elleblythe):

How to solve for √3/2cos50-1/2sin50???

OpenStudy (alekos):

hint: cos30 = sqrt3/2 and sin30 = 1/2

OpenStudy (elleblythe):

@alekos yes from the unit circle right? but how do I solve for cos50 if it isn't on the unit circle??

OpenStudy (alekos):

Works out to cos (30+50) after substitution. So we have cos80 = 0.1736